OpenStudy (anonymous):

What is the value of 12C4? A. 48 B. 3 C. 495 D. 11,880

OpenStudy (anonymous):

@amistre64 Do you think you could help me with this?

OpenStudy (amistre64):

what is 12 P 4?

OpenStudy (amistre64):

12 C 4 = 12 P 4, divided by 4!

OpenStudy (anonymous):

I dont undertand.

OpenStudy (amistre64):

then you most likely have not reviewed your course materials enough

OpenStudy (anonymous):

Its either that or Im tired for doing 4 other lessons, I got 100% on those but I just dont understand that one part wich got me 79% this is my second attempt, ITs just this one stupid problem

OpenStudy (here_to_help15):

12C4 = C(12,4) C(n,r) = n! / [(n - r)! r!] C(12,4) = 12! / [(12 - 4)! r!] C(12,4) = 12! / (8! 4!) C(12,4) = 12 * 11 * 10 * 9 / (4 * 3 * 2) = 990/2 = ?

OpenStudy (amistre64):

how have you tried to work it?

OpenStudy (anonymous):

would I do, 12*11*10 4*3*2*1?

OpenStudy (here_to_help15):

It's easiest if you simplify these first. In the numerator, all but the r highest numbers cancel, leaving you with what I showed. 12 and 4 * 3 canceled leaving you with 11 * 90 = 990 and 2

OpenStudy (amistre64):

pick 4, not 3 12.11.10.9 --------- 4.3.2.1 then simplify yes 12 and 2

OpenStudy (here_to_help15):

@samman8801 do you know how to divide?

OpenStudy (anonymous):

So I get 11880? Then simplify that, And yes

OpenStudy (here_to_help15):

then divide 990/2 is......

OpenStudy (amistre64):

12.11.10.9 --------- 4.3.2.1 11.10.9 ------- 2 11.5.9 it tends to simplify itself out to start with

OpenStudy (amistre64):

11880 = 12 P 4 divide it by 4.3.2.1 = 24

OpenStudy (anonymous):

OHHHHHHHH, Ok I understand this now

OpenStudy (amistre64):

it is rather elementary :) the notation is always a looper tho

OpenStudy (anonymous):

So my answer is 495?